OpenStudy (anonymous):

Another just for fun question , medal for first correct answer with explanation now :) 48÷2(9+3) = ?

OpenStudy (anonymous):

288

OpenStudy (anonymous):

Why would it be that answer? :) I'll give medal to correct answer with how they got that answe

OpenStudy (anonymous):

You do whats inside the paranthesis fisrt which it is 12, then you divide 48 by 2, you get 24

OpenStudy (anonymous):

then multiply 12 and 24 and you get 288

OpenStudy (anonymous):

Well Parenthesis First So 9+3=12 Then 48/2=24 and 24*12=288

OpenStudy (anonymous):

Thanks awesome answers @joselin12 & @So1D :)

OpenStudy (anonymous):

Yw @InstagramDoll no prob :)

OpenStudy (anonymous):

Your Welcome :)